A bullet travelling with a velocity of 16m/s penetrate a wooden block and come to rest in 0.04m find time taken during retardation

Grade:11

1 Answers

Ram Gambhir
6 years ago
Use the third equation of motionAs we are given diplacement=0.04mV that is 0 (as it stops after colliding with wooden block)Initial velocity =16m/s2as=v^2-u^22×4a/100= -16×16a= -3200m/s^2